# Ever-Evolving Memory by Blending and Refining the Past

### 03 Mar 2024 v1

@@ 3.3.3 Ablation Study
Table 4 shows the performance difference between CREEM and its variants excluding refine processes. We evaluate their consistency and memorability against CREEM origin responses in the same way as above. Overall, it was observed that without refining, the responses that are less consistent and has worse memory.
@@
Table 4 shows the performance difference between CREEM and its variants excluding blending right before generating reponse. This setting only gives memories accumulated in the past session, while CREEM’s original method also provides new memories created in the ongoing session until its responding turn. Overall, CREEM consistently produces more consistent and memory-enhanced responses. However, with CC data, CREEM without pre-RG blending is slightly favored in terms of memorability, possibly due to limited retrieval or blend creations. Increasing these hyperparameters is expected to address this issue.
@@ 4 Conclusion
We created an ever-evolving memory scheme to incorporate both current and past memories, actively refining outdated or contradictory information. Despite using a maximum 2-shot LLM prompting approach, in contrast to other fine-tuned methods, CREEM showed superior memory alignment and response generation. Through CREEM, we aim to create a more human-like memory system, fostering further evolution of chat-bot memory systems.
@@ A.2 Blend for first session Prompt
You need to act as a memory expansion model.
Given a set of statements, you should generate 5 high-level insights about {speaker}.
<Statements>:
{text}
<Results about {speaker}>:
@@ A.3 Blend and Refine Prompt
You need to act as a memory expansion model following two steps.
First, given a set of statements, you should generate high-level insights about {speaker}.
Next, with the high-level insights, you should refine the given statements.
Determine whether the new insight is [New], [Redundant], or [Updated] from the given statements. If it is [Redundant] or [Updated], indicate which statements to remove. If it is [New], there are no statements to remove.
Continue these steps twice to make two meaningful insights.
Here are some examples that can be helpful.
<Statements>:
{text}
<Results about {speaker}>:
@@ A.4 Respond Prompt
Appendix B Eval Prompts
B.1 Response Evaluation Prompt
We use ChatGPT to pairwise compare the responses generated by CREEM with the responses generated by $\text{SumMem}_{\text{MSC}}$ and $\textsc{ReBoT}_{\text{CC}}$. In conducting this work, two criteria are used: consistency and memorability. An example of the format is given and the criteria were specified. The criteria according to the evaluation criteria are as follows.
@@ B.1 Response Evaluation Prompt
B.2 Memory Evaluation Prompt
We compare which of the memories created by $\text{SumMem}_{\text{MSC}}$, $\textsc{ReBoT}_{\text{CC}}$, and CREEM had better performance through ChatGPT using the CC dataset and the MSC dataset. There are a total of 7 criteria used in this evaluation, and each prompt contains different evaluation steps for each criteria so that we can evaluate which memory better reflects the conversation by looking at the total dialogue and the memory created for the dialogue. Considering that the results are slightly different each time ChatGPT generates, a total of 20 results are generated at a time and the average is taken as the result. The evaluation step is omitted due to length. The seven criteria are as follows:
  • Conciseness (1-5): This criterion evaluates whether the memory is well-organized, avoiding unnecessary repetition or extraneous information.
@@ Appendix C Examples of Memory construction and Response generation
is her passion, she was an absolute natural with keeping my nerves at bay.
@@
the sun and pool. They also have recommendations for a day visit at the beach."
@@
color that shows when you wear a ponytail."
@@ Appendix F License and Terms
For our implementation and evaluation, we use Huggingface library [^1] https://huggingface.co/ and vLLM library. Both libraries are licensed under Apache License, Version 2.0. We have confirmed that all of the artifacts used in this paper are available for non-commercial scientific use.

### 07 Apr 2024 v2

@@ 5.2 Ablation Study
Table  5 shows excluding the Refine process degrades the response quality. It can be observed that refining leads to better memory retention and consistent responses compared to simply accumulating all information without refinement.
@@
Table 5 shows the performance difference between CREEM and its variants excluding the Insight Infusing process right before generating a response. This setting only gives memories accumulated in the past session, while CREEM’s original method also provides new memories created in the ongoing session until its responding turn. Overall, CREEM consistently produces more consistent and memory-enhanced responses.
Figure 4: Example episode of how past memory is blended and utilized. Change in Speaker A’s hair color choice is reflected in the later sessions.
@@ 6 Conclusion
We create an ever-evolving memory scheme to properly integrate both present and past memories, actively refining outdated or contradictory information. Despite using a maximum 2-shot LLM prompting approach, our method, CREEM, outperforms other fine-tuned methods, $\text{SumMem}_{\text{MSC}}$ and $\textsc{ReBot}_{\text{CC}}$, in terms of memory alignment and response generation. Even when compared with with the $\text{GPT}_{\text{CareCall}}$ which applies different refining methods within the same settings with CREEM using ChatGPT, we have demonstrated the effectiveness of our refining method in improving memory retention and consequently enhancing the quality of responses. Through CREEM, we aim to create a more human-like memory system, fostering further evolution of chat-bot memory systems.
@@ A.2 Blend for first session Prompt
Create insights by extracting detailed information from the ¡Current Dialogue¿.
Here is an example that can be helpful.
[INPUT]
¡Current Dialogue¿
{current_dialogue}
[OUTPUT]
@@ A.3 Blend and Refine Prompt
Create insights by extracting detailed information from the LAST TWO utterences of the ¡Current Dialogue¿ and reflecting the ¡Past Memories¿.
You can make three types of insights. In one insight, include information about only one speaker.
1. [Updated]
If any information from the ¡Past Memories¿ is changed in the ¡Current Dialogue¿, create an Insight reflecting the changes.
Only update the memory with the corresponding speaker’s new information.
Then, indicate which past memory has been changed to remove.
2. [Redundant]
If ¡Current Dialogue¿ contains redundant information of the ¡Past Memories¿, merge them including the details.
Then, indicate which past memory is redundant to remove.
3. [New]
If there is new information that is not related to the ¡Past Memories¿, create a new Insight.
Then, write ”None” and do not remove anything.
Here are some an examples.
[INPUT]
¡Past Memories¿
{retrieved_memories}
¡Current Dialogue¿:
{current_dialogue}
[OUTPUT]
@@ A.4 Respond Prompt
Appendix B Evaluation Prompts
B.1 Memory Evaluation Prompt
We compare which of the memories created by $\text{SumMem}_{\text{MSC}}$, $\textsc{ReBoT}_{\text{CC}}$, $\text{GPT}_{\text{CareCall}}$, and CREEM had better performance through ChatGPT using the CC dataset and the MSC dataset. There are a total of 6 criteria used in this evaluation, and each prompt contains different evaluation steps for each criteria so that we can evaluate which memory better reflects the conversation by looking at the total dialogue and the memory created for the dialogue. Considering that the results are slightly different each time ChatGPT generates, a total of 20 results are generated at a time and the average is taken as the result. The evaluation step is omitted due to length. The seven criteria are as follows:
B.2 Question Generating Prompt
We use ChatGPT to view the entire conversation session and create multiple-choice questions that captured the changes. The prompts used to create the problem are as follows, and examples of prompts are omitted for length: Your task is to create three questions multiple choices that can be answered from the given dialogue. Actively reference the contents of the dialogue in the options, but try to use different expressions.
You will make several questions asking only about the target speaker. The subject of the question should be the target speaker.
Focus on the changed status of the target speaker and use that as a source for making questions.
To make the question difficult, you should follow these steps:
1. Find any information that has changed about the target speaker during the conversation.
The format should be: ”[Before] past information [After] newly updated information”
2. Generate questions that asks about the changed information.
3. Next step is to make multiple answer options that is confusing. Make a correct choice from the newly updated information mentioned after the [After] token, and wrong choices from past outdated information mentioned after the [Before] token.
4. Write the correct answer with a brief explanation.
Here is an example question.
Now generate 3 questions about the target speaker, {target_speaker}.
Target Speaker: {target_speaker}
[Dialogue]
{dialogue}
[Information about {target_speaker}]
B.3 Question Answering Prompt
We use ChatGPT to solve the problems created through the above prompt through the memories of each model. The prompt to solve the problem using memories is as follows: Based on the given memory, solve the three questions by choosing one of the options to answer the following question. Also choose the sentences that support your answer.
Make sure to answer based on ONLY the given memory, without prior knowledge.
The answer might not be directly addressed in the memory so try hard to infer the most reasonable answer among choices A, B, C. But if you still cannot find any answers using the memory, choose ”D) No answer”.
The answer format should be:
[Answers]:
Answer 1:
A) She is a very active and lively.
Supporting sentences:
Sentence 2) She has three kids and her mom is pediatrician.
Sentence 5) She commutes by boat every day to work and lives in amsterdam with canals.
Now solve the questions.
[Memory]:
{memory}
[Questions about Co-worker B]
{questions}
[Answers]:
B.4 Response Evaluation Prompt
We use ChatGPT to pairwise compare the responses generated by CREEM with the responses generated by $\text{SumMem}_{\text{MSC}}$, $\textsc{ReBoT}_{\text{CC}}$ and $\text{GPT}_{\text{CareCall}}$. In conducting this work, two criteria are used: consistency and memorability. An example of the format is given and the criteria were specified. The criteria according to the evaluation criteria are as follows.
@@ Appendix C Examples of Memory construction and Response generation
Speaker A: Oh my youngest did so good! You can tell that being a
cosmetologist is her passion, she was an absolute natural with keeping my nerves at bay.
@@
the sun and pool. They also have recommendations for a day visit at the beach.”
@@
color that shows when you wear a ponytail.
$\text{GPT}_{\text{CareCall}}$
    Memory Speaker A has asked their youngest grandkid to help them dye their hair turquoise,
as they want to get into hairdressing as a profession.
Speaker A jokes that nature will eventually give Speaker B grey hair.
Speaker A recommended the book B̈eautiful Ruinsb̈y Jess Walter to Speaker B
which is a fiction book about Hollywood and actors/actresses.”
Speaker B recently bought a VW Tiguan in a red color.
    Response And speaking of hair, I remembered you just got that red VW Tiguan.
So now we both have a pop of color in our lives!
@@ Appendix F License and Terms
For our implementation and evaluation, we use Huggingface library [^1] https://huggingface.co/. This library licensed under Apache License, Version 2.0. We have confirmed that all of the artifacts used in this paper are available for non-commercial scientific use.